# Partial sampling of a random spanning tree

### v4

@@ Abstract
We present a decomposition of the distribution of the subtree connecting several vertices for spanning trees sampled in a complete graph and a short self-contained derivation of its asymptotic behavior.
00footnotetext: Key words and phrases: Trees, complete graphs00footnotetext: AMS 2020 subject classification: 60C05, 60J10
@@ 1 Introduction
The purpose of this note is to address the following simple question: Given a random spanning tree $\Upsilon$ of the complete graph with $n$ vertices, and a fixed subset $L$ of $d<n$ specific labeled vertices, how are these $d$ vertices connected within the tree when $n$ is large? More precisely, considering the topological type of the smallest subtree $\Upsilon_{L}$ of the random (uniformly sampled) spanning tree connecting them and the graph distances between all leaves and nodes of this subtree, what is their asymptotic distribution as $n$ increases to infinity?
The theorem given below provides an answer in a slightly more general situation: One of the specific vertices denoted by $\Delta$ is chosen to be the root of $\Upsilon$. The weight of a tree is assumed to be proportional to $\kappa^{p_{\Delta}}$, $\kappa$ being a positive parameter and $p_{\Delta}$ the degree of $\Delta$ in the tree. For $\kappa\neq 1$, the distribution we get on spanning tree is not entirely uniform.
@@
However it turns out that the choice of $\kappa$ is irrelevant to answer the question: We show that asymptotically, the topology of $\Upsilon_{L}$ is uniformly chosen among binary trees whose leaves are $L$, and that the graph distances between all leaves and nodes, normalized by $\sqrt{n}$ , converge towards an explicit distribution. As $d$ increases, these distributions (necessarily) form a consistent system.
The question of constructing $\Upsilon_{L}$ can be raised of course on any graph and we show in section 2 how Wilson’s algorithm allows to solve it in full generality. Once objects have been rigorously defined and notations established, the convergence proof given in section 3 has the merit to be short and simple.
As far as we know, this result did not appear as such in the literature. However, it is very close of a description of the Continuous Random Tree given by Aldous in [3]. The consistent system of asymptotic distributions we get can certainly be interpreted in terms of sampling of the CRT. The point here is to show their relevance at the discrete elementary level. More details and references are given at the end of the last section.
2 Definitions, first properties and statement
@@
Consider a graph $\mathcal{G}$ equipped with conductances and a non-vanishing killing function defined on the set of vertices, here denoted by $X$. These data allow to define a Green matrix $G$ indexed by pairs of vertices. Adding a cemetery point $\Delta$, we can consider the value of the killing function at any vertex $x$ as a conductance between $x$ and $\Delta$. Then a spanning forest of rooted trees on $\mathcal{G}$ can be identified to a spanning tree on the extended graph $\mathcal{G}_{\Delta}$, by connecting $\Delta$ to the roots. An extension of Cayley’s theorem (cf. for example [7], section 8-2 in [8]) shows that if we define the weight of such a spanning tree to be the product of the conductances of its edges, the sum of these weights is the determinant of $G$. It provides naturally a probability $\mathbb{P}$ on spanning trees rooted in $\Delta$ and consequently a probability on spanning forests.
@@
This applies in particular to the spanning forests of the complete graph $K_{n}$ with vertices $\{1,2,...,n\}$ endowed with unit conductances and a constant killing function $\kappa$. The corresponding probability on rooted trees will be denoted by $\mathbb{P}^{(n),\kappa}$. If $\kappa=1$, the spanning tree rooted in $\Delta$ can be identified with a uniform spanning tree on $K_{n+1}$. Local limits for these objects have been determined in [4], extending a result of Grimmett [5].
@@
where $J$ denotes the $(n,n)$ matrix with all entries equal to $1$. It is easy to check that for any $(m,m)$ square matrix $M$ with diagonal entries equal to $a+b$ and off diagonal entries equal to $b$, $\det(M)=a^{m-1}(a+mb)$. Hence we get the following identity:
@@
Consequently, in the case of the complete graph with $n$ vertices $\{1,...,n\}$ endowed with unit conductances and a constant killing factor $\kappa$ if $Y$ has $a$ vertices (root excepted) and $r$ edges incident to $\Delta$, by formula 2, we have:
Lemma 2.1
$\mathbb{P}^{(n),\kappa}(Y=\Upsilon_{L})=\frac{\kappa^{r-1}(a+\kappa)}{(n+\kappa)^{d}}$
@@
We say that a tree imbedded in the complete graph and rooted in $\Delta$ is a $L$-tree if it contains $L$ and his set of leaves, denoted by $L^{*}$, is contained in $L$. $\Upsilon_{L}$ is always a $L$-tree. We denote by $N$ the set of nodes of $\Upsilon_{L}$. There is a mapping $j$ from $N\cup L$ into parts of $L$. It maps a vertex $x$ to the set of elements of $L$ which are above it in the tree. In particular, it sends a leaf $i$ into $\{i\}$. We also set $j(\Delta)=L$.
We say that a tree is $L$-reduced if all its vertices which are not elements of $L$ or nodes have been removed. A $L$-tree can be decomposed into its $L$-reduction which is still a $L$ -tree and finite sequences of intermediate inner vertices with a single child.
It is clearly enough to consider the case $L=\{1,...l\}$. Then the natural order on $L$ is used to induce an order on subsets of $L$ by listing the leaves in each set in increasing order and then use alphabetical order to compare their lists of increasing digits. For example we have $(1)<(1,2)<(1,3,7)<(2)<(2,3)$ etc.. Nodes and leaves of a reduced $L$-tree inherit of that order we will call the $L$-order.
In order to capture the way elements of $L$ are connected, in other words the geometry of the subtree connecting them, only the labels of $L$ do matter. This leads us to define an equivalence relation: we say that two $L$-trees are equivalent if they can be exchanged by a permutation of $\{l+1,...,n\}$. Note that equivalence preserves the $L$-order we defined on $N\cup L$. The (equivalence) class of the $L$-reduction of $\Upsilon_{L}$ denoted by $Q_{L}$ inherits a tree structure. Its inner vertices are the images of $N\cup L\setminus L^{*}$ by $j$ and its leaves singletons in $L^{*}$. The class of $\Upsilon_{L}$ is determined by $Q_{L}$ and the numbers of intermediate inner vertices between each vertex in $N\cup L$ and the first vertex in $N\cup L\setminus L^{*}\cup\{\Delta\}$ below it, hence by a map $u_{L}$ from the set of vertices of $Q_{L}$ into the set $\mathbb{N}$ of nonnegative integers. We can say that $(Q_{L},u_{L})$ is a $\mathbb{N}$-extension of $Q_{L}$.
@@
We can now formulate precisely our problem which is to show that as $n$ increases to infinity, the probability that the reduction of $\Upsilon_{L}$ is binary converges to one, and to determine the asymptotic behavior of the joint distribution of the pair $(Q_{L},u_{L})$.
The answer is given in the following:
Theorem 2.1
@@ 3 Proof
Then we start with the proof of c). Note that setting $\Sigma=\sum_{1}^{2l-1}[t_{i}\sqrt{n}]$ there are $\prod_{i=l}^{\Sigma+l}(n-i)$ possible choices for the inner vertices of $\Upsilon_{L}$, given that $u_{L}=[t\sqrt{n}]$. From lemma 2.1, we get that
$\mathbb{P}^{(n),\kappa}(Q_{L}=\alpha,u_{L}=[t\sqrt{n}])=\frac{\Sigma+l+\kappa}{(n+\kappa)^{\Sigma+l}}\prod_{i=l}^{\Sigma+l}(n-i).$
Using Stirling’s approximation, it follows that
@@
$\ln(\mathbb{P}^{(n),\kappa}(Q_{L}=\alpha,u_{L}=[t\sqrt{n}]))=\ln(\Sigma+l+\kappa)-(\Sigma+l)\ln(n+\kappa)+(n-l)\ln(n-l)-n+l\\ +\frac{1}{2}\ln(2\pi(n-l))-(n-\Sigma-l)\ln(n-\Sigma-l)+n-\Sigma-l-\frac{1}{2}\ln(2\pi(n-\Sigma-l))+O(1/n)$ (3)
After decomposing the first term into $1/2\ln(n)+\ln(\Sigma/\sqrt{n})+\ln(1+(l+\kappa)/\Sigma)$ and each term of the form $\ln(n+a)$ into $\ln(n)+\ln(1+a/n)$, Taylor’s formula applied to the $\ln$ functions gives:
@@
$\ln(\mathbb{P}^{(n),\kappa}(Q_{L}=\alpha,u_{L}(i)=[t_{i}\sqrt{n}])=1/2\ln(n)+\ln(\Sigma/\sqrt{n})-l\ln(n)-\Sigma\\ +n\ln(1-\frac{l}{n})-n\ln(1-\frac{\Sigma+l}{n})+O(n^{-1/2})=\ln(\Sigma/\sqrt{n})-(2l-1)\ln(\sqrt{n})-\Sigma^{2}/2n+O(n^{-1/2}).$ (4)
The statement c) follows. One notes that if the $t_{i}$’s are bounded, the remainder term $O(1/n)$ in the first equation can be bounded uniformly by a constant multiple of $1/n$. Similarly, in the last equation, $O(n^{-1/2})$ can be uniformly bounded by a constant multiple of $1/\sqrt{n}$.
To prove b), we can first check that $\int_{\mathbb{R}_{+}^{2l-1}}(\sum_{1}^{2l-1}t_{i})e^{-(\sum_{1}^{2l-1}t_{i})^{2}/2}=c_{l}^{-1}$. Then the following Riemann sum argument shows that b) follows from c): the remark ending the proof of c) shows that given $k\in\mathbb{N}^{2l-1}$ such that $k_{i}<c\sqrt{n}$, $\lim_{n\uparrow\infty}n^{(2l-1)/2}\mathbb{P}^{(n),\kappa}(Q_{L}=\alpha,u_{L}=k)=(\sum_{1}^{2l-1}k_{i}/\sqrt{n})e^{-(\sum_{1}^{2l-1}k_{i})^{2}/2n}$ uniformly in $k$.
Hence $\mathbb{P}^{(n),\kappa}(Q_{L}=\alpha)>\lim_{n\uparrow\infty}\sum_{k\in\mathbb{N}^{2l-1},k_{i}<c\sqrt{n}}\mathbb{P}^{(n),\kappa}(Q_{L}=\alpha,u_{L}=k)=n^{-(2l-1)/2}\sum_{k\in\mathbb{N}^{2l-1},k_{i}<c\sqrt{n}}(\sum_{1}^{2l-1}k_{i}/\sqrt{n})e^{-(\sum_{1}^{2l-1}k_{i})^{2}/2n}+O(n^{-1/2})$˙It converges towards $\int_{[0,c]^{2l-1}}(\sum_{1}^{2l-1}t_{i})e^{-(\sum_{1}^{2l-1}t_{i})^{2}/2}\prod_{i}dt_{i}$ which can be made arbitrarily close from $c_{l}^{-1}$ by taking $c$ large enough. This concludes the proof.
4 Additional comments

### v5

# Biased spanning forests and Gibbs partitions

@@ Abstract
We study biased spanning forests on the complete graph $K_{N}$, obtained by assigning weight proportional to $\kappa^{r}$ to spanning trees on $K_{N}\cup\{\Delta\}$, where $r$ is the degree of a distinguished root $\Delta$.
Fixing a finite set $L$ of marked vertices, we analyse the minimal $\Delta$-rooted subtree connecting $L$ to $\Delta$. In this sense, we investigate the effect of partially sampling a large random spanning forest through finitely many vertices.
For fixed $\kappa$, the reduced subtree is asymptotically a uniformly distributed binary tree, and graph distances rescaled by $\sqrt{N}$ converge jointly in distribution to an explicit limit introduced by Aldous in his study of the Brownian CRT (continuum random tree).
We show that the scale $\kappa\asymp\sqrt{N}$ is critical: if $\kappa=o(\sqrt{N})$, the marked vertices lie asymptotically in a single component; if $\kappa\gg\sqrt{N}$, the induced partition is asymptotically discrete and the distances to $\Delta$ are negligible on the $\sqrt{N}$ scale.
In the critical regime $\kappa=c\sqrt{N}$, the induced partition of the marked vertices converges to a non-degenerate limit law. This limit is an explicit Gibbs partition of type $\alpha=\tfrac{1}{2}$, whose coefficients admit a closed one-dimensional integral representation. We also note that the limiting EPPF we obtain coincides with a conditional excursion-partition law for reflected Brownian bridges identified by Pitman.
Our approach remains entirely discrete and combinatorial: all asymptotics are obtained from explicit determinant formulas and elementary expansions, without invoking continuum limits.
@@ 1 Introduction
Given a random spanning tree $\Upsilon$ of the complete graph $K_{N}$ with $N$ vertices, and a fixed subset $L=\{a_{1},\dots,a_{l}\}$ of $l<N$ labelled vertices, a natural question is the following: how are these $l$ vertices connected within the tree when $N$ is large? More precisely, considering the smallest subtree $\Upsilon_{L}$ of $\Upsilon$ containing $L$, what is the asymptotic distribution, as $N\to\infty$, of its topological type and of the graph distances between its leaves and internal nodes?
This problem can be viewed as a finite-dimensional description of the local geometry of large random spanning trees.
Spanning forests with killing.
We formulate the problem in a slightly more general setting. We distinguish one vertex $\Delta$ as a root and consider random spanning trees on $K_{N}\cup\{\Delta\}$ with weight proportional to $\kappa^{r}$, where $\kappa>0$ is a parameter and $r$ denotes the degree of $\Delta$. Equivalently, when restricted to $K_{N}$, this induces a $\kappa$-biased spanning forest in which a forest with $r$ components has weight proportional to $\kappa^{\,r-1}$.
When $\kappa$ is fixed, this model differs only mildly from the uniform spanning tree. However, when $\kappa$ varies with $N$, qualitatively new phenomena appear. In particular, the scale $\kappa\asymp\sqrt{N}$ turns out to be critical. Our approach remains entirely discrete and combinatorial: all asymptotics are obtained from explicit determinant formulas and elementary expansions, without invoking continuum limits.
Main mechanism.
The asymptotic analysis throughout the paper follows a single principle. For fixed reduced tree shapes and fixed integer edge-length vectors, the Green-determinant formula yields an explicit finite-$N$ expression. After a suitable $\sqrt{N}$ rescaling of the edge lengths, this expression admits a limit. Summing these limits over all integer length vectors and all binary reduced shapes produces Riemann sums which converge to explicit integrals. A crucial point is that the sum of these integrals equal $1$. This shows that asymptotically no probability mass is lost outside the $\sqrt{N}$ scaling window and outside the class of binary reduced trees. Binary shapes therefore emerge with total limiting mass $1$, and their distribution becomes uniform by symmetry.
Fixed-$\kappa$ regime.
@@
In Section 2 we show that, for fixed $\kappa$ and fixed $m$, the value of $\kappa$ is asymptotically irrelevant for the local structure of $\Upsilon_{L}$: the reduced tree of $\Upsilon_{L}$ is asymptotically uniformly distributed among binary trees with leaves $L$, and the graph distances, rescaled by $\sqrt{N}$, converge jointly to an explicit distribution. As $l$ varies, these distributions form a consistent family.
This family coincides with the consistent family of “proper $k$-trees” introduced by Aldous in his study of the Brownian CRT (continuum random tree). In contrast, the present paper derives these asymptotic laws purely by discrete enumeration and elementary asymptotics, without appealing to continuum constructions.
Critical regime and phase transition.
A different phenomenon occurs when
$\kappa=c\sqrt{N},\qquad c>0.$
This scale is critical. If $\kappa=o(\sqrt{N})$, the induced partition of $L$ is asymptotically trivial: with high probability all marked vertices lie in a single component. If $\kappa\gg\sqrt{N}$, the opposite triviality holds: the induced partition is asymptotically the discrete partition, and the graph distance from each marked vertex to $\Delta$ is $o(\sqrt{N})$. Only in the critical regime $\kappa\sim c\sqrt{N}$ do these two effects balance, producing a non-degenerate limiting partition.
In Section 3 we show that in this regime the reduced object $Q_{L}$ obtained from $\Upsilon_{L}$ is, with high probability, a bouquet consisting of $r\in\{1,\dots,l\}$ binary trees attached to $\Delta$. Moreover, for each fixed $r$ and each fixed bouquet configuration (i.e. a partition $\pi$ of $L$ into $r$ blocks, together with a binary tree on each block), all such configurations have the same limiting probability, namely
$\operatorname{\mathbb{P}}\bigl(Q_{L}=\text{given bouquet with $r$ trees}\bigr)\longrightarrow I_{l,r}(c),$
[…]
$\operatorname{\mathbb{P}}\bigl(|\Pi_{N}|=r\bigr)\longrightarrow C_{l,r}\,I_{l,r}(c),$
where $|\Pi_{N}|$ is the number of blocks of the induced partition of $L$ and $C_{l,r}$ is the number of binary forests on $l$ labelled leaves with $r$ components. Conditionally on $|\Pi_{N}|=r$, the internal subtrees defined by the blocks remain asymptotically independent, following the same binary-tree limit law as in the fixed-$\kappa$ case.
Explicit Gibbs structure.
We show in section 3 that the limiting partition in the regime $\kappa\sim c\sqrt{N}$ is Gibbs of type $\alpha=\tfrac{1}{2}$ in the sense of Gnedin and Pitman [8], with block weights $w_{m}=(2m-3)!!$ and explicitly computable coefficients $V_{l,r}(c)$. These weights are normalized by the identity $\sum_{r=1}^{l}C_{l,r}\,I_{l,r}(c)=1$, proved in Section 3. This family is not of Poisson–Dirichlet type, and arises here from a concrete combinatorial model. It appears that the limiting EPPF in the critical regime agrees with Pitman’s conditional excursion partition for a reflected Brownian bridge given its local time at 0.
Dyck paths and excursion partitions.
This latter fact is clearly related to the existence of a path-based encoding of the tree $\Upsilon_{L}$. Once an order is fixed on the leaves (for instance the LL–order used throughout), the $\Delta$–rooted tree $\Upsilon_{L}$ admits a canonical contour/height exploration, whose height process is a Dyck path of length twice the number of edges. Under this encoding, the blocks of the induced partition of the marked vertices correspond to excursion intervals of the height process.
Multitype extension
A multi-cemetery extension can be treated similarly: Wilson’s algorithm can be run in such a way that whenever a loop-erased walk is killed, one records which cemetery point was used. This produces a spanning forest whose components are coloured by cemetery labels, but we focus on the single-root case as ”multitype” versions of the construction and of the asymptotics study require no new ideas.
The relation with previous work is discussed in the final section.
2 Spanning trees in the complete graph
2.1 Definitions, first properties and statement
Green determinants.
@@ 2.1 Definitions, first properties and statement
Consider a graph $\mathcal{G}$ equipped with conductances and a non-vanishing killing function defined on the set of vertices, here denoted by $X$. These data allow to define a Green matrix $G$ indexed by pairs of vertices. Adding a cemetery point $\Delta$, we can consider the value of the killing function at any vertex $x$ as a conductance between $x$ and $\Delta$. Then a spanning forest of rooted trees on $\mathcal{G}$ can be identified to A spanning tree on the extended graph $\mathcal{G}_{\Delta}$, rooted in $\Delta$ induces a spanning forest on $\mathcal{G}$. An extension of Cayley’s theorem (cf. for example [11], section 8-2 in [12]) shows that if we define the weight of such a spanning tree to be the product of the conductances of its edges, the sum of these weights is the inverse of the determinant of $G$. It provides naturally a probability $\mathbb{P}$ on spanning trees rooted in $\Delta$ (and consequently a probability on spanning forests.)
This applies in particular to the complete graph $K_{N}$ with vertices $\{1,2,...,N\}$ endowed with unit conductances and a constant killing function $\kappa$. It is naturally associated with a random walk for which at each step, the random walk is killed with probability $\frac{\kappa}{\kappa+N-1}$.
@@
The corresponding probability on $\Delta$-rooted spanning trees will be denoted by $\mathbb{P}^{(N),\kappa}$. If $\kappa=1$, the spanning tree rooted in $\Delta$ can be identified with a uniform spanning tree on $K_{N+1}$. Local limits for these objects have been determined in [6], extending a result of Grimmett [9].
@@
where $J$ denotes the $(N,N)$ matrix with all entries equal to $1$. It is easy to check that for any $(d,d)$ square matrix $M$ with diagonal entries equal to $a+b$ and off diagonal entries equal to $b$, $\det(M)=a^{d-1}(a+db)$. Hence we get the following identity:
@@
Consequently, in the case of the complete graph with $N$ vertices $\{1,...,N\}$ endowed with unit conductances and a constant killing factor $\kappa$, if $Y$ has $d$ vertices (root excepted) and $r$ edges incident to $\Delta$, by formula 2, we have:
Lemma 2.1.
$\mathbb{P}^{(N),\kappa}(Y=\Upsilon_{L})=\frac{\kappa^{r-1}(d+\kappa)}{(N+\kappa)^{d}}$
Reduced trees and lengths.
@@
We say that a tree imbedded in the complete graph and rooted in $\Delta$ is a $L$-tree if it contains $L$ and his set of leaves, denoted by $L^{*}$, is contained in $L$. $\Upsilon_{L}$ is always a $L$-tree. We denote by $\mathcal{N}$ the set of nodes of $\Upsilon_{L}$.
There is a mapping $j$ from $\mathcal{N}\cup L$ into parts of $L$. It maps a vertex $x$ to the set of elements of $L$ which are above it in the tree. In particular, it sends a leaf $i$ into $\{i\}$. We also set $j(\Delta)=L$.
We say that a tree rooted in $\Delta$ is $L$-reduced if all its vertices which are not elements of $L$ or nodes have been removed. A $L$-tree can be decomposed into its $L$-reduction which is still a $L$ -tree and finite sequences of intermediate inner vertices with a single child.
It is clearly enough to consider the case $L=\{1,...l\}$. Then the natural order on $L$ is used to define an order on the children of each internal node: we order these children by the minimum element of the set of leaves below them. (In the binary case this simply fixes a canonical left/right orientation at each branch point.) This turns $Y$ into a plane rooted tree in a deterministic way.
Perform the depth-first contour traversal of the resulting plane rooted tree, starting and ending at $\Delta$; this visits each edge exactly twice and defines an order on all vertices of $Y$.
In order to capture the way elements of $L$ are connected, in other words the geometry of the subtree connecting them, only the labels of $L$ do matter. This leads us to define an equivalence relation: we say that two $L$-trees are equivalent if they can be exchanged by a permutation of $\{l+1,...,N\}$. Note that equivalence preserves the $L$-order we defined on $\mathcal{N}\cup L$. The (equivalence) class of the $L$-reduction of $\Upsilon_{L}$ denoted by $Q_{L}$ inherits a tree structure. Its inner vertices are the images of $\mathcal{N}\cup L\setminus L^{*}$ by $j$ and its leaves singletons in $L^{*}$. The class of $\Upsilon_{L}$ is determined by $Q_{L}$ and the numbers of intermediate inner vertices between each vertex in $\mathcal{N}\cup L$ and the first vertex in $\mathcal{N}\cup L\setminus L^{*}\cup\{\Delta\}$ below it, hence by a map $u_{L}$ from the set of vertices of $Q_{L}$ into the set $\mathbb{N}$ of nonnegative integers. We can say that $(Q_{L},u_{L})$ is a $\mathbb{N}$-extension of $Q_{L}$.
@@
Contour (Dyck-path) encoding.
We recall a standard encoding of rooted trees by walks on $\operatorname{\mathbb{Z}}_{\geq 0}$. A Dyck path of length $2n$ is a function $H:\{0,1,\dots,2n\}\to\operatorname{\mathbb{Z}}_{\geq 0}$ such that $H(0)=H(2n)=0$ and $H(k+1)-H(k)\in\{+1,-1\}$ for all $k$.
Let $v_{k}$ be the visited vertex at step $k$ by the contour path. If $Y$ is a rooted tree with $n$ vertices, denoting by $d_{Y}$ the graph distance in $Y$, set
$H_{Y}(k):=d_{Y}(\Delta,v_{k}),\qquad k=0,1,\dots,2n,$
Then $H_{Y}$ is a Dyck path of length $2n$.
[…]
Asymptotics.
We can now formulate precisely our problem which is to show that as $N$ increases to infinity, the probability that the reduction of $\Upsilon_{L}$ is binary converges to one, and to determine the asymptotic behavior of the joint distribution of the pair $(Q_{L},u_{L})$. This is given in the following:
Theorem 2.1.
@@
i.e. the distribution of $Q_{L}$ converges towards the uniform distribution on $\mathcal{B}_{l}$.
@@ 2.2 Proof
Then we start with the proof of c). Note that setting $\Sigma=\sum_{1}^{2l-1}[t_{i}\sqrt{N}]$ there are $\prod_{i=0}^{\Sigma-1}(N-l-i)$ possible choices for the inner vertices of $\Upsilon_{L}$, given that $u_{L}=[t\sqrt{N}]$. From lemma 2.1, we get that
$\mathbb{P}^{(N),\kappa}(Q_{L}=\alpha,u_{L}=[t\sqrt{N}])=\frac{\Sigma+l+\kappa}{(N+\kappa)^{\Sigma+l}}\prod_{i=0}^{\Sigma-1}(N-l-i).$
(with the usual convention that an empty product equals 1).
We define
$L_{N}:=\log(\frac{\Sigma+l+\kappa}{(N+\kappa)^{\Sigma+l}}\prod_{i=0}^{\Sigma-1}(N-l-i)),$
set $\Sigma=\sigma\sqrt{N}$, and assume that $0\leq\sigma\leq\sigma_{\max}$, with $l$ and $\kappa$ fixed.
We prove that there exist constants $C>0$ and $N_{0}$ (depending only on $\sigma_{\max},l,\kappa$) such that for all $N\geq N_{0}$ and all $\sigma\in[0,\sigma_{\max}]$,
@@
$\left|\,L_{N}-\Bigl(\log(\Sigma+l+\kappa)-l\log N-\frac{\sigma^{2}}{2}\Bigr)\right|\leq\frac{C}{\sqrt{N}}.$
Uniform estimate for the critical scaling
Recall
$L_{N}=\log(\Sigma+l+\kappa)-(\Sigma+l)\log(N+\kappa)+\sum_{i=0}^{\Sigma-1}\log(N-l-i),$
with $\Sigma=\sum_{i=1}^{2l-1}k_{i}$ and $\Sigma\leq\sigma_{\max}\sqrt{N}$.
Rewrite
$\log(N+\kappa)=\log N+\log\!\left(1+\frac{\kappa}{N}\right),\qquad\log(N-l-i)=\log N+\log\!\left(1-\frac{l+i}{N}\right),$
so that
$L_{N}=\log(\Sigma+l+\kappa)-l\log N-(\Sigma+l)\log\!\left(1+\frac{\kappa}{N}\right)+\sum_{i=0}^{\Sigma-1}\log\!\left(1-\frac{l+i}{N}\right).$
Choose $N_{0}$ such that for all $N\geq N_{0}$ and $\Sigma\leq\sigma_{\max}\sqrt{N}$,
$0\leq\frac{\kappa}{N}\leq\frac{1}{2},\qquad 0\leq\frac{l+\Sigma-1}{N}\leq\frac{1}{2}.$
(1) The term $\log(1+\kappa/N)$. For $0\leq y\leq 1/2$,
$0\leq\log(1+y)\leq y+y^{2}.$
Hence
$\left|(\Sigma+l)\log\!\left(1+\frac{\kappa}{N}\right)\right|\leq(\Sigma+l)\left(\frac{\kappa}{N}+\frac{\kappa^{2}}{N^{2}}\right).$
Using $\Sigma\leq\sigma_{\max}\sqrt{N}$,
$\left|(\Sigma+l)\log\!\left(1+\frac{\kappa}{N}\right)\right|\leq\frac{\kappa(\sigma_{\max}+l)}{\sqrt{N}}+\frac{\kappa^{2}(\sigma_{\max}+l)}{N^{3/2}}\leq\frac{C_{1}}{\sqrt{N}},$
where $C_{1}=\kappa(\sigma_{\max}+l)+\kappa^{2}(\sigma_{\max}+l)$.
(2) The product term. For $0\leq x\leq 1/2$,
$\log(1-x)=-x+\varepsilon(x),\qquad|\varepsilon(x)|\leq\frac{x^{2}}{1-x}\leq 2x^{2}.$
Set $x_{i}=(l+i)/N$. Then
$\sum_{i=0}^{\Sigma-1}\log\!\left(1-\frac{l+i}{N}\right)=-\frac{1}{N}\sum_{i=0}^{\Sigma-1}(l+i)+\sum_{i=0}^{\Sigma-1}\varepsilon(x_{i}).$
Moreover,
$\left|\sum_{i=0}^{\Sigma-1}\varepsilon(x_{i})\right|\leq 2\sum_{i=0}^{\Sigma-1}\left(\frac{l+i}{N}\right)^{2}\leq\frac{2}{N^{2}}\,\Sigma\,(l+\Sigma)^{2}\leq\frac{C_{2}}{\sqrt{N}},$
with $C_{2}=2\sigma_{\max}(l+\sigma_{\max})^{2}$.
Now
$\sum_{i=0}^{\Sigma-1}(l+i)=\Sigma l+\frac{\Sigma(\Sigma-1)}{2},$
so
$\frac{1}{N}\sum_{i=0}^{\Sigma-1}(l+i)=\frac{\Sigma^{2}}{2N}+\frac{\Sigma l}{N}-\frac{\Sigma}{2N}.$
Hence
$\left|\frac{1}{N}\sum_{i=0}^{\Sigma-1}(l+i)-\left(\frac{\Sigma^{2}}{2N}+\frac{\Sigma l}{N}\right)\right|=\frac{\Sigma}{2N}\leq\frac{\sigma_{\max}}{2\sqrt{N}}=:\frac{C_{3}}{\sqrt{N}}.$
Combining,
$\sum_{i=0}^{\Sigma-1}\log\!\left(1-\frac{l+i}{N}\right)=-\frac{\Sigma^{2}}{2N}-\frac{\Sigma l}{N}+r_{N},\qquad|r_{N}|\leq\frac{C_{2}+C_{3}}{\sqrt{N}}.$
(3) Conclusion. Inserting (1) and (2) into $L_{N}$ yields
$\left|L_{N}-\left(\log(\Sigma+l+\kappa)-l\log N-\frac{\Sigma^{2}}{2N}\right)\right|\leq\frac{C}{\sqrt{N}},$
with $C=C_{1}+C_{2}+C_{3}+l\,\sigma_{\max}$, uniformly for $\Sigma\leq\sigma_{\max}\sqrt{N}$.
Write
@@
$\delta_{N}:=L_{N}-\Bigl(\log(\Sigma+l+\kappa)-l\log N-\frac{\Sigma^{2}}{2N}\Bigr),\qquad|\delta_{N}|\leq\frac{C}{\sqrt{N}}.$
Then
$e^{L_{N}}=(\Sigma+l+\kappa)\,N^{-l}e^{-\Sigma^{2}/(2N)}\,e^{\delta_{N}}.$
For $N$ large enough so that $C/\sqrt{N}\leq 1$, we have $|e^{\delta_{N}}-1|\leq 2|\delta_{N}|$, hence
$\left|e^{L_{N}}-(\Sigma+l+\kappa)\,N^{-l}e^{-\Sigma^{2}/(2N)}\right|\leq\frac{2C}{\sqrt{N}}\,(\Sigma+l+\kappa)\,N^{-l}e^{-\Sigma^{2}/(2N)},$
uniformly for all $\sigma\in[0,\sigma_{\max}]$. This yields the uniform estimate in c). Moreover, for all $N$ sufficiently large and all $k\in\mathbb{N}^{2l-1}$ satisfying $k_{i}\leq\sigma_{\max}\sqrt{N}$,
$\left|N^{(2l-1)/2}\mathbb{P}^{(N),\kappa}(Q_{L}=\alpha,u_{L}=k)-\left(\sum_{i=1}^{2l-1}\frac{k_{i}}{\sqrt{N}}\right)\exp\!\left(-\frac{(\sum_{i=1}^{2l-1}k_{i})^{2}}{2N}\right)\right|\leq\frac{C}{\sqrt{N}}.$
To prove b), we first compute the normalising integral:
$\int_{\mathbb{R}_{+}^{2l-1}}\left(\sum_{i=1}^{2l-1}t_{i}\right)\exp\!\left(-\frac{(\sum_{i=1}^{2l-1}t_{i})^{2}}{2}\right)dt=c_{l}^{-1}.$
Fix $\sigma_{\max}>0$. By the uniform estimate obtained in part (c),
$\liminf_{N\to\infty}\mathbb{P}^{(N),\kappa}(Q_{L}=\alpha)$ $\geq\liminf_{N\to\infty}\sum_{\begin{subarray}{c}k\in\mathbb{N}^{2l-1}\\ k_{i}\leq\sigma_{\max}\sqrt{N}\end{subarray}}\mathbb{P}^{(N),\kappa}(Q_{L}=\alpha,u_{L}=k)$
$=\lim_{N\to\infty}N^{-(2l-1)/2}\sum_{\begin{subarray}{c}k\in\mathbb{N}^{2l-1}\\ k_{i}\leq\sigma_{\max}\sqrt{N}\end{subarray}}\left(\sum_{i=1}^{2l-1}\frac{k_{i}}{\sqrt{N}}\right)\exp\!\left(-\frac{(\sum_{i=1}^{2l-1}k_{i})^{2}}{2N}\right).$
The last sum is precisely the Riemann sum over the cube $0,\sigma_{\max}]^{2l-1}$ with mesh size $h_{N}=1/\sqrt{N}$ of the continuous function
$f(t)=\left(\sum_{i=1}^{2l-1}t_{i}\right)\exp\!\left(-\frac{(\sum_{i=1}^{2l-1}t_{i})^{2}}{2}\right).$
Therefore,
$\liminf_{N\to\infty}\mathbb{P}^{(N),\kappa}(Q_{L}=\alpha)\geq\int_{[0,\sigma_{\max}]^{2l-1}}f(t)\,dt.$
Letting $\sigma_{\max}\to\infty$, monotone convergence gives
$\liminf_{N\to\infty}\mathbb{P}^{(N),\kappa}(Q_{L}=\alpha)\geq\int_{\mathbb{R}_{+}^{2l-1}}f(t)\,dt=c_{l}^{-1}.$
Since the probabilities $\mathbb{P}^{(N),\kappa}(Q_{L}=\alpha)$ sum to $1$ over the $c_{l}$ binary shapes, we must have equality:
$\lim_{N\to\infty}\mathbb{P}^{(N),\kappa}(Q_{L}=\alpha)=c_{l}^{-1}.$
This completes the proof of b).
2.3 Additional comments